what is the exact value of: tan25 +tan110/ 1-tan25 tan110
Answer by lwsshak3(11628)

what is the exact value of: tan25 +tan110/ 1-tan25 tan110
use tan addition formula:
tan(s+t)=(tan s+tan t)/(1-tan s tan t)
For given problem:
(tan25 +tan110)/ (1-tan25 tan110)
=tan(25+110)
=tan (135)
=-1 (ref angle of 45º in quadrant II where tan<0)